# From Logits to Hierarchies: Hierarchical Clustering made Simple

### v1

@@ 4 Experiments
In this section we showcase the experimental results obtained with our proposed approach. In the first part, we focus on the task of hierarchical clustering. We demonstrate that existing specialized models for hierarchical clustering present major limitations when applied in realistic settings. In contrast, our proposed approach achieves convincing results in challenging vision datasets, markedly outperforming alternative methods. In the second part of this section, we present a case study where we discuss the application of the L2H algorithm on top of a pre-trained ImageNet classifier, demonstrating its value for model interpretability and the discovery of spurious correlations.
@@ 4.1 Hierarchical clustering
In this section, we compare the performance of our proposed method for hierarchical clustering with recent specialized approaches on three challenging vision datasets: namely the CIFAR-10, CIFAR-100 (Lake et al., 2015) and Food-101 (Bossard et al., 2014) datasets. We report our results in Table 1. For each dataset, we implement our algorithm on top of two pre-trained flat clustering models, namely TURTLE (Gadetsky et al., 2024) and TEMI (Adaloglou et al., 2023). These are two state-of-the-art clustering methods (see Section B.2 for more details), both of which are not designed to produce a hierarchy of clusters. In our evaluation, we report both metrics to evaluate models at the flat level and metrics to evaluate the quality of the produced hierarchy. For comparing models at the flat level, we report NMI, ARI, Accuracy and Leaf Purity (LP). To assess the quality of the hierarchical clustering, we report two metrics: Dendrogram Purity (DP) and Least Hierarchical Distance (LHD). The former was introduced in Kobren et al. (2017) and extends the notion of purity, normally evaluated at the leaf level, to assess the quality of a tree clustering: higher purity corresponds to higher quality of the hierarchy. Note that this metric was recently adopted in Manduchi et al. (2023) to benchmark deep hierarchical clustering models. Least Hierarchical Distance, on the other hand, measures the average minimal log-distance in the hierarchy between any pair of datapoints that have the same true label but different cluster assignments. A better hierarchy corresponds to a lower LHD. More details about our metrics can be found in the Appendix Section B.3.
The results in Table 1 uncover the aforementioned limitations of recent deep learning methods (DeepECT, TreeVAE). In particular, these models do not scale well in terms of the depth of the hierarchy, thereby producing overly shallow hierarchies for datasets with a large number of classes (CIFAR-100, Food101). We find this to be linked to their high computational complexity, which we discuss later in Table 2. For instance, TreeVAE learns a hierarchical generative model with leaf-specific decoders: this choice helps its performance in a generative scenario but impacts its scalability to large-scale datasets. Importantly, the comparison in terms of flat clustering metrics highlights that ad-hoc hierarchical models produce clusterings at the leaf level that are much less accurate than those obtained with non-hierarchical models. The results for DP and LHD highlight the shortcomings in these tasks. Note that we notice the presence of artifacts, underlined in Table 1, for the LHD metric, due to overly shallow hierarchies that degenerate to having only two to three leaves.
@@
NMI ()($\uparrow$)( ↑ ) ARI ()($\uparrow$)( ↑ ) ACC ()($\uparrow$)( ↑ ) LP ()($\uparrow$)( ↑ ) DP ()($\uparrow$)( ↑ ) LHD ()($\downarrow$)( ↓ )
CIFAR-10
Agglomerative 0.074 0.038 0.211 0.246 0.121 0.549 10
@@
L2H-TEMI 0.778 0.565 0.682 0.698 0.502 0.298 100
L2H-Turtle 0.917 0.831 0.896 0.896 0.803 0.235 100
Food-101
Agglomerative 0.082 0.004 0.039 0.045 0.011 0.438 101
@@
Table 1: Quantitative comparison of hierarchical clustering performance on three datasets (CIFAR-10, CIFAR-100, Food-101). We report as a baseline Agglomerative clustering (Agg), deep hierarchical specialized models (DeepECT, TreeVAE), and our L2H method applied on top of two state-of-the-art flat models (TEMI, TURTLE). We also indicate the number of leaves in the hiearchy modelled by each approach, and whether a given method can perform inference on a hold-out test set. We bold best results for each metric and underline results that are artifacts of degenerate solutions with shallow hierarchies. Results are averaged over five runs.
@@
In contrast to alternative approaches, our proposed algorithm recovers high-quality hierarchies for all three datasets, when implemented on top of both TEMI and TURTLE models. Looking at both hierarchical metrics, our method markedly outperforms other approaches, with a consistent margin over costly deep learning specialized methods. These findings demonstrate that L2H can leverage the information embedded in the logits of a pre-trained flat clustering model to model an accurate hierarchy of the clusters, outperforming ad-hoc hierarchical models, even without having access to internal representations. Note as well that our method does not require any fine-tuning of the pre-trained model and, by construction, retains its clustering performance at the leaf level, which matches state-of-the-art in our results. Finally, the efficacy of our proposed procedure is not hindered by the presence of a large number of the classes in the dataset, as we witness for other methods.
@@
Figure 2: Visualization of the hierarchical clustering produced by L2H-TURTLE on the CIFAR-100 dataset. The inferred hierarchy is represented as a circular tree. On the lowest level, the leaves are annotated by reporting the most frequent label for the samples in each leaf. Leaves are color-colored according to the 20 superclasses in the dataset.
In practice, hierarchical clustering results are often used as a visualization tool, and to analyze the structure of a dataset at different levels of granularity. Hence, to evaluate our proposed approach we visualize and inspect the hierarchy obtained with L2H-TURTLE on the CIFAR-100 dataset in Figure 2. Note that given the absence of leaf labels, we associate a class label to each leaf by looking at the most frequent label among the datapoints in the given leaf. While an off-the-shelf ground-truth hierarchy is not available for the CIFAR-100 dataset, the authors organize the 100 classes in 20 superclasses. Hence, we color-code the inferred leaf labels in the hierarchy by superclasses and check if the hierarchical clustering recovers this global structure. Notably, the global structure of the superclasses is largely reflected in the visualized hierarchy. Most interesting is that the outliers, for which the color does not coincide with the neighboring leaves, still reflect meaningful semantic associations. For instance, whale and dolphin—despite being aquatic mammals—are grouped with fish species. However, this is not surprising, given their adaptation exclusively to aquatic environments and the presence of similar traits to fishes, like streamlined bodies. On the contrary, mammals such as otter, beaver, and seal, which are only semi-aquatic, are grouped with other small to medium-sized terrestrial mammals, emphasizing size and communal characteristics like the presence of limbs and fur. Another example is the characterization of worm and snake alongside in the hierarchy. Although snakes are reptiles, their elongated, limbless bodies visually resemble those of non-insect invertebrates like worms. This showcased analysis confirms the efficacy of our method in recovering a tree structure that follows meaningful semantic associations. The results indicate that our method produces hierarchies that enable detailed exploration of the structure in the data at varying levels of granularity. Inspecting the hierarchy gives valuable insights for interpretability, revealing underlying associations by the model.
Finally, we end this section with a comparison in terms of the computational cost of our method compared with alternative models, and in particular with specialized deep learning approaches for hierarchical clustering. As the results in Table 2 show, our proposed L2H algorithm is extremely lightweight. For completeness, we also measure the overall runtime to perform hierarchical clustering with our method, including the runtime to train the TURTLE model, as an example model used to compute logits as input. Our approach allows to perform hierarchical clustering extremely efficiently even on large-scale datasets such as ImageNet-1K, with a total training time of a few minutes. Note that, due to the combined efficiency of our method and state-of-the-art flat clustering models, the overall runtime scales seamlessly with dataset size and number of leaves in the hierarchy. Conversely, alternative hierarchical deep learning approaches exhibit a significantly higher computational cost. Moreover, a large dataset size and number of classes markedly increase the computational burden.
@@
L2H $<0.01$ $<0.01$ $<0.01$ $0.47$
Agglomerative $0.1$ $0.1$ $0.9$
HypHC $163.7$ $153.3$ $195.3$
DeepECT $24.1$ $26.2$ $67.5$
TreeVAE $364.1$ $756.3$ $2293.7$
L2H-TURTLE $1.6$ $1.6$ $1.7$ $5.27$
@@
Table 2: Training time (in minutes) for our proposed method compared to baselines for hierarchical clustering on CIFAR-10, CIFAR-100, and Food-101 datasets. At the top, we report the runtime for the L2H algorithm alone. Below, we report the runtime of the TURTLE model plus our L2H algorithm to produce a hierarchy, compared with the runtime of each baseline model. Results are averaged over three runs.
4.2 Case study: Pre-trained ImageNet Classifier
Next, we demonstrate how our method can be used in a supervised setup to produce a hierarchical clustering given the logits of a pre-trained classifier. Specifically, we use the ImageNet-1K dataset Deng et al. (2009), which comprises over a million images and a thousand distinct classes with an underlying hierarchical structure. We apply L2H on the logits of a pre-trained ImageNet classifier to determine the hierarchy of classes. As a pre-trained classifier, we use the InternImage model (Wang et al., 2022). The resulting hierarchical clustering is visualized in Figure 3.
Figure 3(a) shows the inferred tree for the thousand ImageNet classes. The colors indicate whether a leaf node corresponds to the superclass “artifact” or “organism”, which we determine based on the corresponding WordNet hypernyms of each class. Overall, we observe a distinct separation between the two superclasses in the inferred tree. Figure 3(b) zooms into a subtree of the inferred tree that comprises different bird species. Specifically, it shows 58 of the 60 classes of birds found in the ImageNet dataset. The leaf nodes are colored by different clades of bird species (based on the WordNet hierarchy), showing that the inferred hierarchy groups together related species. For example, the group “aquatic bird” is almost completely represented in one of the two main branches, which further splits into a separate cluster for “parrots” and another one for “bird of prey”. The other main branch of the tree subdivides further into “passerine” and “game bird” forming distinct clusters.
Overall, our results suggest that the inferred hierarchy recovers a significant portion of the global and local hierarchical structure of the ImageNet dataset given the logits of a pre-trained ImageNet classifier trained with non-hierarchical labels. Yet, the inferred tree also reveals interesting outliers. For example, in Figure 3(a), there is a distinct subtree for snow-related artifacts (e.g., dogsled, snowmobile, bobsled) within a large branch of the tree that comprises organisms. Further investigation shows that this group of artifacts is merged with arctic animals (e.g., malamute, Siberian husky, Eskimo dog), which reveals an intuitive correlation between classes and highlights potential biases of the pre-trained model. Likewise, in Figure 3(b), we see potential outliers such as “peacock” among the group of parrots or “bustard” among game birds. While the WordNet hierarchy classifies “bustard” as a wading bird, hence among aquatic birds, the inferred hierarchy places it among game birds. Interestingly, common definitions of bustards as terrestrial game birds support the inferred categorization.
@@ 4.2 Case study: Pre-trained ImageNet Classifier
Figure 3: Visualization of the hierarchical clustering produced by L2H-TURTLE for the ImageNet-1K dataset. The inferred hierarchy is represented as a circular tree, where the leaf nodes are organized in a circle. Figure 3(a) shows the complete tree colored by the corresponding WordNet hypernyms “artifact” and “organism”, which are the largest two superclasses in the ImageNet dataset. Figure 3(b) shows the subtree of birds colored by different bird species if they comprise more than one class. The results show that our method recovers a significant portion of the global and local hierarchical structure of the ImageNet dataset.
@@ 5 Conclusion
In this work, we propose a lightweight yet effective procedure for hierarchical clustering based on pre-trained non-hierarchical models. Notably, our solution proves to be markedly more effective and significantly more computationally efficient than alternative methods. Different from existing models for hierarchical clustering, our method can successfully handle large datasets of many classes, taking an important step in deploying hierarchical clustering methods in challenging settings. Moreover, we show that the usefulness of our approach extends to supervised setups, by implementing it on top of a pre-trained classifier to recover a meaningful hierarchy of classes. A case study on ImageNet shows that this approach provides relevant insights for interpretability, and can reveal potential biases of the pre-trained model and ambiguities in existing categorizations.
While we provide extensive results on image datasets, in this work we do not explore other data modalities. However, our procedure is general and may be applied to different data types, which we leave for future work. Hierarchical clustering presents important advantages over non-hierarchical clustering by simultaneously capturing the structure in the data at multiple levels of granularity. However, inspecting the hierarchy is still necessary to extract valuable insights. An important focus for future work is to investigate strategies to partly bypass this process, automatically selecting levels of the hierarchies that provide the most meaningful clustering.
Reproducibility Statement
We share a Python implementation of our L2H algorithm both in Appendix A and in the supplementary material. To ensure the reproducibility of our experimental results, in Appendix B we provide detailed insights on datasets and metrics, as well as on the implementation of our method and the compared baselines.

### v2

@@ 4 Experiments
In this section, we present the experimental results of our study. In the first part, we empirically demonstrate that existing specialized deep hierarchical clustering models face significant limitations in realistic scenarios. These limitations stem from their high computational demands and their difficulty in handling a large number of clusters. In contrast, our proposed method demonstrates compelling results on challenging vision datasets, achieving substantially better performance compared to these specialized models.
While the first part of this section focuses on hierarchical clustering, the second part extends our approach to supervised setups. We provide a case study showcasing the application of the L2H algorithm on top of a pre-trained ImageNet classifier. The results illustrate its potential for enhancing model interpretability and uncovering spurious correlations. Moreover, they complement the results for hierarchical clustering, demonstrating that our method can also be applied on top of pre-trained supervised models. Further details on datasets, implementations and metrics can be found in Appendix B.
@@ 4.1 Hierarchical Clustering
In this section, we evaluate the performance of recent specialized hierarchical clustering approaches on three challenging vision datasets— CIFAR-10, CIFAR-100 (cifarpaper) and Food-101 (Bossard:2014aa)— comparing their performance with our proposed method. We show our results in Table 1, including metrics to evaluate models at the leaf level and metrics to evaluate the quality of the produced hierarchy. To compare model performance at the leaf level, we report Normalized Mutual Information (NMI), Adjusted Rand Index (ARI), Accuracy and Leaf Purity (LP). To assess the quality of the hierarchical clustering, we report two metrics: Dendrogram Purity (DP) and Least Hierarchical Distance (LHD). The former was introduced in dppaper and extends the notion of purity, normally evaluated at the leaf level, to assess the quality of a tree clustering: higher purity corresponds to higher quality of the hierarchy. Note that this metric was recently adopted in manduchi2023tree to benchmark deep hierarchical clustering models. Least Hierarchical Distance, on the other hand, measures the average minimal log-distance in the hierarchy between any pair of data points that have the same true label but different cluster assignments. A better hierarchy corresponds to a lower LHD. More details about our metrics can be found in Section B.3. For each dataset, we implement our L2H algorithm on top of two pre-trained flat clustering models, namely TURTLE (gadetsky2024letlabelsunsupervisedtransfer) and TEMI (Adaloglou2023TEMI). These are two recent clustering methods (see Section B.2 for more details), both of which are not designed to produce a hierarchy of clusters but only a flat clustering.
@@
NMI $(\uparrow)$ ARI $(\uparrow)$ ACC $(\uparrow)$ LP $(\uparrow)$ DP $(\uparrow)$ LHD $(\downarrow)$
CIFAR-10 Agglomerative 0.074 0.038 0.211 0.246 0.121 0.549 10
@@
L2H-TEMI 0.778 0.565 0.682 0.701 0.502 0.298 100
L2H-Turtle 0.917 0.831 0.896 0.897 0.803 0.235 100
Food-101 Agglomerative 0.082 0.004 0.039 0.045 0.011 0.438 101
@@
Table 1: Quantitative comparison of hierarchical clustering performance on three datasets (CIFAR-10, CIFAR-100, Food-101). We report as a baseline agglomerative clustering, deep hierarchical specialized models (DeepECT, TreeVAE), and our L2H method applied on top of two state-of-the-art flat models (TEMI, TURTLE). We also indicate the number of leaves in the hiearchy modelled by each approach, and whether a given method can perform inference on a hold-out test set. We bold best results for each metric and underline results that are artifacts of degenerate solutions with shallow hierarchies. Notably the application of L2H does not affect flat clustering performance, retaining the clustering performance of the pre-trained model (TURTLE, TEMI) at the leaf level.
The results in Table 1 uncover the aforementioned limitations of recent deep hierarchical clustering methods (DeepECT, TreeVAE), which fail to achieve satisfactory performance even on moderately challenging datasets such as CIFAR-10. The results prove that these approaches struggle at modeling deep hierarchies, producing overly shallow trees in datasets with a large number of classes such as CIFAR-100 or Food-101. This is also linked to their high computational complexity. For instance, TreeVAE learns a hierarchical generative model with leaf-specific decoders: this choice helps its performance in a generative scenario but impacts its scalability to large-scale datasets (see Table 2 for more details). Moreover, the comparison in terms of flat clustering metrics highlights that deep hierarchical models produce clusterings at the leaf level that are much less accurate than those obtained with non-hierarchical models. [^3] In constrast, our approach retains the strong performance of the backbone model (TURTLE or TEMI) at the leaf level.
@@
In contrast to alternative approaches, our proposed method recovers high-quality hierarchies for all three datasets when implemented with both TEMI or TURTLE as the backbone model. The results in the hierarchical metrics demonstrate that our method markedly outperforms existing approaches, with a consistent margin over costly deep learning specialized models. These findings demonstrate that the L2H algorithm can leverage the information embedded in the logits of a pre-trained flat clustering model to model an accurate hierarchy of the clusters. Most importantly, they show that our approach outperforms sophisticated deep hierarchical models while being more scalable and efficient. Note as well that our method does not require any fine-tuning of the pre-trained model, nor access to the internal representations. By construction, it retains the clustering performance of the pre-trained model at the leaf level, which matches state-of-the-art in our results. Importantly, the efficacy of our method is not hindered by the presence of a large number of classes in the dataset, as we witness for other methods. In particular, in Appendix C we show that our method achieves remarkable hierarchical clustering results on datasets as large as ImageNet-1k (imagenetpaper). As well we show in Appendix C that our approach is applicable across different choices for the backbone model, and notably, it outperforms deep specialized hierarchical approaches even when the chosen backbone model yields weaker clustering performance at the leaf level than TURTLE or TEMI. We also report additional ablations, and in particular we validate the robustness of our approach with respect to the hyperparameter $K$, corresponding to the number of clusters at the leaf level of the hierarchy.
@@
Figure 2: Visualization of the hierarchical clustering produced by L2H-TURTLE on the CIFAR-100 dataset. The inferred hierarchy is represented as a circular tree. On the lowest level, the leaves are annotated by reporting the most frequent label for the samples in each leaf. Leaves are color-coded according to the 20 superclasses in the dataset.
@@
L2H <0.01<0.01 <0.01<0.01 <0.01<0.01 $0.45\pm 0.0$
Agglomerative <0.1<0.1 <0.1<0.1 $0.8$ -
HypHC $163.7\pm 4.0$ $153.3\pm 19.4$ $195.3\pm 3.2$ -
DeepECT $24.1\pm 15.1$ $26.2\pm 9.8$ $67.5\pm 36.6$ -
TreeVAE $364.1\pm 76.8$ $756.3\pm 178.6$ $2293.7\pm 211.7$ -
L2H-TURTLE $1.6\pm 0.0$ $1.6\pm 0.0$ $1.7\pm 0.0$ $5.25\pm 0.0$
@@
Table 2: Training time (in minutes) for our proposed method compared to baselines for hierarchical clustering on CIFAR-10, CIFAR-100, and Food-101 datasets. At the top, we report the runtime for the L2H algorithm alone. Below, we report the runtime of the TURTLE model plus our L2H algorithm to produce a hierarchy, compared with the runtime of each baseline model. Results are averaged over three runs and include standard deviations.
In practice, hierarchical clustering results are often used as a visualization tool and to analyze the structure of a dataset at different levels of granularity. Hence, to evaluate our proposed approach, we visualize and inspect the hierarchy obtained with L2H-TURTLE on the CIFAR-100 dataset in Figure 2. Note that given the absence of leaf labels, we associate a class label to each leaf by looking at the most frequent label among the data points in the given leaf. While an off-the-shelf ground-truth hierarchy is not available for the CIFAR-100 dataset, the authors organize the 100 classes in 20 superclasses. Hence, we color-code the inferred leaf labels in the hierarchy by superclasses and check if the hierarchical clustering recovers this global structure. Notably, the global structure of the superclasses is largely reflected in the visualized hierarchy. Most interesting is that the outliers, for which the color does not coincide with the neighboring leaves, still reflect meaningful semantic associations. For instance, whale and dolphin—despite being aquatic mammals—are grouped with fish species. However, this is not surprising, given their adaptation exclusively to aquatic environments and the presence of similar traits to fishes, like streamlined bodies. On the contrary, mammals such as otter, beaver, and seal, which are only semi-aquatic, are grouped with other small to medium-sized terrestrial mammals, emphasizing size and communal characteristics like the presence of limbs and fur. Another example is the characterization of worm and snake alongside in the hierarchy. Although snakes are reptiles, their elongated, limbless bodies visually resemble those of non-insect invertebrates like worms. This showcased analysis confirms the efficacy of our method in recovering a tree structure that follows meaningful semantic associations. The results indicate that our method produces hierarchies that enable detailed exploration of the structure in the data at varying levels of granularity. Inspecting the hierarchy gives valuable insights for interpretability, revealing underlying associations by the model.
We end this section with a comparison in terms of the computational cost of our method compared with alternative models, and in particular with specialized deep learning approaches for hierarchical clustering. As the results in Table 2 show, our proposed L2H algorithm is extremely lightweight. To compare with the efficiency of alternative methods, we also measure the overall runtime to perform hierarchical clustering with L2H-TURTLE, which includes the runtime to train the TURTLE model, as an example of backbone model. Our approach allows us to perform hierarchical clustering extremely efficiently, even on large-scale datasets such as ImageNet-1k, with a total training time of a few minutes. Note that, due to the combined efficiency of our method and state-of-the-art flat clustering models like TURTLE, the overall runtime scales seamlessly with dataset size and number of leaves in the hierarchy. Conversely, deep hierarchical approaches exhibit a significantly higher computational cost. Moreover, as with TreeVAE, increasing dataset size and number of classes markedly increases the computational burden. Finally, it is to note that TURTLE, as well as TEMI, leverages CLIP embeddings for training. Hence, to further validate the superior efficiency and efficacy of our method compared to alternative deep specialized approaches, in Appendix C we compare our results with the ones obtained by training TreeVAE using CLIP embeddings. These results again validate that our L2H approach achieves markedly superior performance when compared models have access to foundation model embeddings for training. Once again, this gain in performance comes also with much higher computational efficiency.
@@
Figure 3: Visualization of inferred hierarchy for the ImageNet-1K dataset. The hierarchy is represented as a circular tree, where the leaf nodes are organized in a circle. Figure 3(a) shows the complete tree colored by the corresponding WordNet hypernyms “artifact” and “organism”, which are the largest two superclasses in the ImageNet dataset. Figure 3(b) shows the subtree of birds colored by different bird species if they comprise more than one class. The results show that our method recovers a significant portion of the global and local hierarchical structure of the ImageNet dataset.
4.2 Case Study: Pre-trained ImageNet Classifier
In this section we complement the results from the previous section by showing that our method can be applied in a supervised setup, producing a hierarchy given the logits of a pre-trained classifier. Specifically, we use the ImageNet-1k dataset (imagenetpaper), which comprises over a million images and a thousand distinct classes, with an underlying hierarchical structure organized according to the WordNet hierarchy. We apply the L2H algorithm on the logits of a pre-trained ImageNet classifier to model the hierarchy of classes. As a pre-trained classifier, we use the InternImage model (wang2022internimage).
Figure 3(a) illustrates the inferred hierarchy for the thousand ImageNet classes. The colors indicate whether a leaf node corresponds to the superclass “artifact” or “organism”, which are the largest two superclasses that can be determined based on the corresponding WordNet hypernyms of each class in ImageNet. Overall, we observe a distinct separation between the two superclasses in the inferred tree.
In addition, Figure 3(b) shows a subtree of the inferred hierarchy that comprises different bird species. Specifically, it shows 58 of the 60 classes of birds contained in the ImageNet dataset. The leaf nodes are colored by different clades of bird species (based on the WordNet hierarchy), showing that the inferred hierarchy groups together related species. For example, the group “aquatic bird” is almost completely represented in one of the two main branches, which further splits into a separate cluster for “parrots” and another one for “bird of prey”. The other main branch of the tree subdivides further into “passerine” and “game bird” forming distinct clusters.
Overall, our results suggest that the inferred hierarchy recovers a significant portion of the global and local hierarchical structure of the ImageNet dataset given the logits of a pre-trained ImageNet classifier trained with non-hierarchical labels. Yet, the inferred tree also reveals interesting outliers. For example, in Figure 3(a), there is a distinct subtree for snow-related artifacts (e.g., dogsled, snowmobile, bobsled) within a large branch of the tree that comprises organisms. Further investigation shows that this group of artifacts is merged with arctic animals (e.g., malamute, Siberian husky, Eskimo dog), which reveals a spurious correlation between classes and highlights potential biases of the pre-trained model. Likewise, in Figure 3(b), we see potential outliers such as “bustard” among game birds. Interestingly, this constitutes an example where our method reveals an ambiguity in the WordNet hierarchy, which classifies “bustard” as a wading bird, while it is usually defined as a terrestrial game bird, like in our inferred categorization.
@@ 5 Conclusion
In this work, we uncover significant limitations of existing deep hierarchical clustering models, demonstrating that these methods face important challenges in scaling to large-scale datasets, falling short of delivering satisfactory performance. As an alternative approach, we propose a lightweight yet effective procedure for hierarchical clustering based on pre-trained non-hierarchical models. Notably, our solution proves to be markedly more effective and significantly more computationally efficient than existing methods. Specifically, we demonstrate that our method can successfully handle large datasets with hundreds of classes, taking an important step for the practical applicability of hierarchical clustering methods in realistic settings. Moreover, we show that the usefulness of our approach extends to supervised setups by implementing it on top of a pre-trained classifier to recover a meaningful hierarchy of classes. A case study on ImageNet shows that our approach provides relevant insights for interpretability, and can reveal potential biases in the pre-trained model or spurious correlation in the data.
Our proposed method is general and may be applied to different data types beyond vision, which we leave as an opportunity for future work. Another direction for future work is the investigation of strategies for automatically selecting meaningful levels of the inferred hierarchy. Hierarchical clustering presents important advantages over non-hierarchical clustering by simultaneously capturing the structure in the data at multiple levels of granularity. However, manually inspecting the hierarchy is still necessary to extract valuable insights. Thus, future work could investigate strategies to partly bypass this process, automatically selecting levels of the hierarchy that provide the most meaningful clustering. Our proposed method relies on the assumption that logits from a pre-trained flat model can be used as a proxy to measure cluster similarities. While empirically supported by our experimental results, this assumption may not always hold, for instance if the pre-trained model is poorly calibrated.
Acknowledgements
EP is supported by a fellowship from the ETH AI Center, and received funding from the grant #2021-911 of the Strategic Focal Area “Personalized Health and Related Technologies (PHRT)” of the ETH Domain (Swiss Federal Institutes of Technology). AR is supported by the StimuLoop grant #1-007811-002 and the Vontobel Foundation. The authors are grateful to Laura Manduchi for useful discussions.
Impact Statement
As other ML models, clustering methods—such as the ones explored in this work—are vulnerable to the risk of reflecting potential biases and spurious correlations from the data they are trained on. However, hierarchical approaches offer a promising direction for mitigating these risks by enhancing the interpretability and transparency of clustering and classification outcomes.